Question 301 pts

When using the allowance method to estimate bad debt, what will an entry to write off an account receivable using the allowance method do?

Group of answer choices

increase net income

decrease net income

have no effect on net income

increase liabilities

Question 311 pts

On January 1, our company purchased a truck for $80,000. The estimated useful life of the truck is 4 years. The residual value at the end of 4 years is estimated to be $10,000. What is the depreciation expense for the third year of use if we use the straight-line method?

Group of answer choices

$17,500

$20,000

$35,000

$52,500

Question 323 pts

Use the following information to answer these questions.

On January 1, our company purchased a truck for $80,000. The estimated useful life of the truck is 4 years. The residual value at the end of 4 years is estimated to be $10,000.

What is the depreciation expense for the second year of use if we use the double-declining balance method?

                           [ Select ]                       ["$40,000", "$35,000", "$22,500", "$20,000"]      

What is the balance in accumulated depreciation at the end of the second year of use if we use the double-declining balance method?

                           [ Select ]                       ["$40,000", "$35,000", "$60,000", "$57,500"]      

What is the book value at the end of the second year of use if we use the double-declining balance method?

                           [ Select ]                       ["$20,000", "$22,500", "$45,000", "$40,000"]      

Question 33

On January 1 of the current year (Year 1), our company acquired a truck for $75,000. The estimated useful life of the truck is 5 years or 100,000 miles. The residual value at the end of 5 years is estimated to be $5,000. The actual mileage for the truck was 22,000 miles in Year 1 and 27,000 miles in Year 2. What is the book value at the end of the first year of use (Year 1) after we complete the adjusting entry if we use the units of production method?

Group of answer choices

$59,600

$56,100

$54,600

$51,100
Answer #1

The answer has been presented in the supporting sheet. For detailed answer refer to the supporting sheet.

A C D Answer Part 1) The correct answer is have no effect on net income Explanation Under allowance method when ever accountA Part 3) B C D E F G H The correct answer is $ 20000 Calculation Double declining rate = 1/4*100*2 = 50% Depreciation for YeA B C D E C) The correct answer is $ 20000 Explanation WDV =Cost - accumulated depreictaion upto year 2 = 80000 - 60000 = = $